I would like to compare what I have to what you get please....

Chapter 2 Problems:

(2.1) An investor recently purchased a corporate bond that yields 7.68%. The investor is in the 25% federal-plus-state tax bracket. What is the bonds after-tax yield to the investor?

(2-2) Corporate bonds issued by Johnson Corporation currently yield 8.0%. Municipal bonds of equal risk currently yield 5.5%. At what personal tax rate would an investor be indifferent between these two bonds?

(2-5) Kendall Corners Inc. recently reported net income of $3.1 million and depreciation of $500,000. What was its net cash flow? Assume it had no amortization expense.

(2-11) The Wendt Corporation reported $50 million of taxable income. Its federal tax rate was 21% (ignore any possible state corporate taxes)

  1. What is the company’s federal income tax bill for the year?
  2. Assume the firm receives an additional $1 millions of interest income from some bonds it owns. What is the additional tax on this interest income?
  3. Now assume that Wendt does not receive the interest income but does receive and additional $1 million as dividends on some stock it owns. Recall that 50% of dividends received are tax exempt. What is the additional tax on this dividend income?

(2-12) The Shrieves Corporation has %10,000 that it plans to invest in marketable securities. It is choosing among AT&T bonds (which yield 6.6%), AT&T preferred stock (with a dividend yield of 6.0%), and state of Florida muni bonds (which yield 5% but are not taxable). The federal tax rate is 21% (ignore any possible state corporation taxes). Recall that 50% of dividends received are tax exempt. Find the after-tax rates of return on all three securities after paying federal corporate taxes.
Answer #1

2.1

After tax rate = YTM * (1-Tax rate)
After tax rate = 7.68 * (1-0.25)
After tax rate = 5.76

2.2

Municipal bond rate = Taxed bond rate * (1-Tax rate)
5.5 = 8 * (1-Tax rate)
Tax rate% = 31.25

2-5

Cash flow = depr+net income = 3100000+500000= 3600000